Two forces of magnitude P and 3^1/2P acthe at right angles to each other.Their resultant makes an angle theta with P.which of the following gives the correct value of theta

Grade:11

1 Answers

J Anto Pradeep
5 years ago
Let us take the resultant vector to be C.
Then the resultant vector C can be written as
 \vec{C} = P\hat{i} + \sqrt{3}P \hat{j}
 
 
tan\theta =\frac{\sqrt{3}P}{P}
tan\theta =\sqrt{3}
 
Then
\theta =60^{\circ}
 
Name: J Anto Pradeep
Contact No.: 9600741702